# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ V-J Machine Learning
For example, the authors in [215] address the minimization of energy and time consumption in mobile-edge computing-enabled HAPS networks. It formulates an optimization problem for user association, service sequencing, and task allocation, and proposes a SVM-FL algorithm to predict user associations without sharing private data. Simulations demonstrate up to 15.4% improvement in energy and time efficiency compared to terrestrial centralized methods. To enable global-scale FL in NTNs, a distributed hierarchical FL framework is proposed [216], HAPS as intermediate servers. Integrating LEO satellites and ground clients while using GEO and MEO satellites for model relaying, the framework enhances privacy, improves model accuracy, manages latency, and boosts scalability. Numerical results confirm its efficiency in training loss reduction and network optimization. Addressing the limitations of traditional ML in satellite communications, the study [217] introduces a federated learning framework that employs HAPS as distributed parameter servers for LEO satellite constellations. Proposed framework integrates a hierarchical communication architecture along with model dissemination and aggregation algorithms, significantly reducing training time from days to hours while improving model accuracy.
VI Open Research Challenges
The integration of TNs and NTNs through HAPS-enabled connectivity represents a transformative opportunity for future 6G wireless communication systems. However, this paradigm presents several challenges, including balancing the trade-off between coverage and capacity, optimizing radio resource allocation, developing efficient beamforming techniques, addressing mobility and handover issues, and ensuring seamless interaction between TN and NTN components. Moreover, the deployment of HAPS systems in dense urban environments requires careful planning to avoid interference and ensure efficient spectrum utilization, particularly when multiple HAPS and ground stations are involved. In the following, we state the hottest research challenges:
VI-1 Channel Modeling
HAPS in NTNs introduce unique channel modeling challenges due to their stratospheric operation (17–25 km altitude) and highly dynamic environment. Unlike satellites or terrestrial infrastructure, HAPS experience mixed terrestrial and space-like propagation conditions with high variability. Key challenges include channel non-stationarity, caused by platform movement, altitude shifts, and atmospheric changes, which impact Doppler shift, delay spread, and shadowing; and channel consistency, ensuring validity across wide coverage areas with diverse terrain and weather. Environmental effects—rain, clouds, atmospheric gases—become significant at mmWave and higher frequencies, leading to frequency-selective fading and attenuation. Moreover, antenna array impairments from platform instability complicate beamforming and degrade link reliability.
VI-2 Spectrum
Since HAPS systems share frequencies with satellites and terrestrial systems, spectrum allocation and interference management are also important issues. To prevent interference, particularly in highly-occupied frequency bands, HAPS networks must effectively coordinate their spectrum utilization. Efficient spectrum sharing and interference mitigation techniques, such as advanced beamforming, interference cancellation, and cognitive radio methods, must be developed to ensure reliable communication without signal degradation.
VI-3 Mobility management
The high mobility of NTN components demands dynamic 3D positioning and load balancing to optimize resource distribution, handovers, and power use. Continuity and latency pose additional challenges when HAPS link both satellite and terrestrial networks, making seamless, low-latency handovers difficult for time-sensitive applications like AR, autonomous driving, and industrial automation. Addressing these issues requires advances in multi-connectivity protocols and intelligent routing to adapt paths based on network state, traffic, and UE needs. Improved synchronization techniques are also essential to reduce latency spikes, and ensure reliable, responsive performance in integrated TN-NTN systems.
VI-4 Interoperability
Seamless integration of TNs and NTNs into a unified 6G network architecture requires overcoming significant standardization and interoperability challenges. Currently, terrestrial and satellite systems operate via distinct communication standards and protocols. Developed 3GPP standards provide a good base for interoperability, but many features of the protocol stuck requires optimization—or a new design all together— to accommodate for the larger delays and massive number of connected UEs. In short, the inclusion of HAPS as a central component of TN-NTN networks will require new protocols and standardization efforts to ensure compatibility across diverse systems. This includes designing common interfaces and protocols for satellites, HAPS, and terrestrial nodes, as well as ensuring that HAPS can dynamically interact with both space- and ground-based infrastructure in real-time.
VI-5 Energy
Energy efficiency is a major challenge for HAPS in NTNs, especially for future 6G systems operating under strict energy constraints. Sustainable operation requires energy-aware communication protocols and power management strategies, as HAPS rely on limited resources like solar power to support communication, sensing, and navigation. Key advancements include improved energy harvesting, battery storage, dynamic power allocation, and resource optimization. While efficient daytime solar harvesting is vital, storing energy for nighttime or low-sunlight conditions remains difficult, making enhanced storage technologies and advanced scheduling essential to extend mission duration and ensure uninterrupted service.
VI-6 Legal regulations
HAPS will share airspace with manned and unmanned aircraft, as well as space-bound vehicles, which raises regulatory and safety concerns about air traffic control, collision avoidance, and airspace management. The expanding role of HAPS in global communication infrastructure will require the evolution of international aviation and telecommunications regulations to ensure the safe integration of HAPS with existing air traffic.
VI-7 Deployment costs
The large-scale deployment of HAPS hinges on economic feasibility, requiring business models that offset high initial costs through mechanisms such as public–private partnerships, rural connectivity subsidies, or integration into smart city projects. Demonstrating long-term benefits—enhanced coverage, rapid deployment, and reduced infrastructure costs—is essential to justify investment. Overcoming regulatory, safety, and cost-effectiveness challenges will be critical to realizing the full potential of HAPS in future networks.
VII Discussion and conclusion

### v2

@@ VI-J Machine Learning
Recent work shows that federated learning (FL) can strengthen HAPS-enabled networks by enhancing privacy and lowering communication overhead. Approaches like support vector machine (SVM)-based FL for task offloading and hierarchical FL with HAPS as distributed servers improve model accuracy, energy efficiency, and reduce training time for large-scale applications. For example, Wang et al.[223] addressed the minimization of energy and time consumption in mobile-edge computing-enabled HAPS networks. It formulates an optimization problem for user association, service sequencing, and task allocation, and proposes a SVM-FL algorithm to predict user associations without sharing private data. Simulations demonstrate up to 15.4% improvement in energy and time efficiency compared to terrestrial centralized methods. Addressing the limitations of traditional ML in satellite communications, Elmahallawy and Luo [64] introduced a FL framework that employs HAPS as distributed parameter servers for LEO satellite constellations. The proposed framework integrates a hierarchical communication architecture along with model dissemination and aggregation algorithms, significantly reducing training time from days to hours while improving model accuracy. To enable global-scale FL in NTNs, a distributed hierarchical FL framework was proposed by Farajzadeh, Yadav and Yanikomeroglu [66], with HAPS as intermediate servers. Integrating LEO satellites and ground clients while using GEO and MEO satellites for model relaying, the framework enhances privacy, improves model accuracy, manages latency, and boosts scalability. Numerical results confirm its efficiency in training loss reduction and network optimization. Güven and Karabulut Kurt [85] proposed a CNN-aided receiver for HAPSLEO links that jointly addresses channel estimation and carrier frequency offset (CFO) estimation under high Doppler and rapid channel variations. Supervised CNN models are trained to learn the nonlinear mapping from received pilot observations to channel and CFO estimates, reducing estimation errors compared to conventional pilot-based methods. Simulation results demonstrated improved synchronization accuracy and detection performance in representative HAPS–LEO scenarios. Mahyastuty et al. [147] investigated CNN-assisted and CNN-based Successive Interference Cancellation (SIC) replacement for PD-NOMA in HAPS-enabled mMTC and WSN scenarios, where the HAPS acts as an aggregation node. The proposed approach targets low-complexity massive access by leveraging CNNs for multiuser detection, and simulation results showed improved BER performance compared with conventional PD-NOMA receivers under the considered system configurations. Later, Mahyastuty, Pamukti, and Oktaviana [148] introduced a CNN-based decoding framework for power-domain NOMA in HAPS systems to overcome the limitations of successive interference cancellation (SIC), particularly error propagation and sensitivity to user dynamics. The study analyzed the impact of user location and mobility on decoding performance, showing that users closer to the HAPS achieve target BER at lower SNR, while increasing user density raises the required SNR for neighboring users.
Table VI: Mapping between learning paradigms, HAPS-specific constraints, representative applications, and open challenges in HAPS-enabled NTN.
Learning Paradigm HAPS-Specific Constraints Addressed Representative Applications in HAPS/NTN Pending Challenges and Open Issues
DRL / RL [163, 28, 165, 195, 141, 111, 157, 130] Mobility and non-stationarity (platform/user dynamics, intermittent visibility). Real-time decision latency (online control and scheduling). Energy budget for onboard inference/training. Platform positioning and 3D trajectory control (HAPS/UAV). Dynamic RA, power control, and interference mitigation. User scheduling under outdated/partial CSI. Joint beamforming and RIS configuration. Sim-to-real gap; robustness to channel/traffic model mismatch. Sample inefficiency and training instability in non-stationary NTN. Multi-agent scalability and coordination overhead. Safety/constraint satisfaction (hard QoS, interference, energy).
FL [223, 66, 64] Privacy constraints (no raw-data sharing). Limited backhaul / feeder link capacity. Intermittent connectivity and long RTTs in NTN. User association and task offloading prediction at the edge. HAPS-assisted distributed training (HAPS as aggregator/parameter server). Global-scale hierarchical learning across ground–HAPS–satellite tiers. Communication overhead of model updates; compression/quantization needs. Stragglers and synchronization under intermittent links. Statistical heterogeneity (non-IID data) and personalization requirements. Secure aggregation and adversarial/poisoning robustness.
CNN [85, 147, 148] High Doppler, residual CFO, and fast channel variations in HAPS/LEO links. Strong multiuser interference and imperfect SIC in NOMA-based HAPS systems. Non-stationary propagation, traffic dynamics, and domain shift across scenarios. CNN-based channel and CFO estimation for HAPS-assisted and HAPS–LEO links. CNN-based multiuser detection/decoding for PD-NOMA in HAPS and mMTC scenarios. Improved BER, throughput, and robustness compared to conventional receivers. Dependence on representative training datasets; sensitivity to domain mismatch. Increased computational and energy cost at HAPS/UE receivers. Limited interoperability and deployment trade-offs (accuracy vs. latency).
VI-K Cybersecurity and Adversarial Threats in HAPS-Enabled NTN
Cybersecurity threats can put NTN systems at significant risk. Due to exposed wireless links and strong reliance on Positioning, Navigation, and Timing (PNT), platforms such as HAPS are particularly vulnerable, making robust security mechanisms essential. By mapping threats to HAPS subsystems (such as communication payload, power management, telemetry, tracking, and flight control, as well as ground station interfaces), Hjaiji, Ouni and Alouini [95] provided a HAPS-focused cybersecurity analysis. The authors specifically highlighted jamming, replay, and system intrusion attacks, data manipulation and spoofing, DoS/DDoS, adversarial ML attacks, and supply-chain risks; and extensively reviewed the resent advancements in defense mechanisms applicable to HAPS subsystems. A simulation-based distributed-denial-of-service case study that quantifies service/control-plane availability deterioration under assault was also provided as a significant technical component of the study.
Yu et al. [239] examined the communication between several transmitters and a multi-antenna UAV when jamming is present. They suggested cooperative beamforming techniques that reduce deliberate jamming and inter-user interference. The analysis showed how anti-jamming performance is affected by geometry, UAV motion, and beam design. In addition to HAPS-assisted NTN systems, it offers practical techniques for aerial-layer resiliency. Using the minimum variance distortionless response (MVDR) beamforming approach, Xi, Liu and Zhao [230] investigated anti-jamming for satellite ground-station arrays. After creating a signal model that takes direction-dependent array/reflector effects into consideration, proposed approach calculates MVDR weights to reduce interference. Simulations verified the improvements in output SINR in relation to jammer angle, power, and snapshot availability. Aiming array-level interference reduction in SatCom/NTN feeder lines, the study may be potentially generalized to HAPS-related application. Otay, Humadi, and Karabulut Kurt [169] investigates a security-relevant scenario in which a HAPS prevents LEO-to-ground communication by acting as a downlink jammer. A direct satellite-ground link with a jamming HAPS and a cooperative two-satellite path (direct+relay) under the same HAPS jammer are the two scenarios that are examined. The authors derive analytical formulas based on outages that quantify the degradation of link dependability caused by jammer power and shape. One important lesson is that satellite cooperation and relaying can lower the likelihood of a serious disruption because, in the worst scenario, both pathways must be jammed at the same time. Cao et al. [42] adapts spectrum-access choices for LEO satellite downlinks under jamming by using multi-agent deep reinforcement learning. Agents choose channels and actions that minimize the impact of jamming while maintaining throughput by learning policies from observation histories. Performance was compared to less adaptable techniques using a jammer model with channel-switching dynamics. For NTN resource management, including HAPS-based architectures, it offers a practical illustration of autonomy-driven anti-jamming. Gu, Xing, and You [81] proposed a global navigation satellite systems spoofing detection approach based on coupled visual/inertial/GNSS positioning algorithm. To lower false alarms, it incorporates confidence handling and residual/consistency testing in the fused navigation solution. Datasets and experimental settings that mimic spoofing effects are used to validate the method. It is pertinent to aerial platforms where secure control and safe flying depend on strong PNT. A lightweight 1D-CNN detector targeted at small UAVs with an emphasis on early detection of global positioning system (GPS) spoofing was proposed by Sung et al. [205]. It reports practical runtime/energy feasibility on embedded hardware and assesses detection performance against classical ML baselines. The technique supports defensive actions (such as safe landing logic) and was shown under practical UAV operational restrictions. In layered NTN ecosystems, it is a powerful applicable reference for UAVs security.
The issue of huge, nearly simultaneous handovers in NTN (e.g., LEO footprints moving quickly), which might overwhelm control signaling and create security gaps, was addressed by Zhang et al. [242]. It suggests a safe group-handover protocol that combines steps to lower signaling costs while maintaining authorization and authentication objectives. The study compares cost and delay to baseline methods and assesses performance using discrete-event simulation. For LEO-to-device links, Coppola et al. [51] develop a physical-layer ranging technique that can withstand spoofing and relay-style manipulation. It examines attack surfaces and presents a safe range waveform/processing chain that works with realistic OFDM-style systems. The authors assess ranging accuracy and resilience under representative satellite channel effects and offer security justifications. The end product is a tangible foundation for NTN systems’ secure timing/ranging (PNT-adjacent) services. Wu et al. [228] examine covert communication in an integrated HAPSTN, approaching security from a low likelihood of detection perspective. To conceal the signal from a warden or eavesdropper, it employs a constant-power auxiliary node that sends artificial noise. The authors present closed-form formulas for outage probability and effective covert rate after deriving the covert constraint connecting transmitter and auxiliary node powers. These findings apply to HAPS deployments where a design objective under surveillance is to conceal transmission behavior (not just encryption).
VII Open Research Challenges
A key direction for future 6G systems is the integration of TNs and NTNs through HAPS-enabled connectivity. Although several related challenges have been partially studied in terrestrial, satellite, and standalone HAPS networks, many issues remain unresolved—or become more critical—under large-scale, multi-tier TN/NTN integration. In particular, heterogeneous propagation conditions, wide coverage areas, platform mobility, and stringent latency and reliability constraints introduce new cross-layer coupling effects spanning the physical, MAC, and network layers. The following summarizes the key open research challenges (see Table VII), clearly distinguishing HAPS-specific open issues from those already identified in prior work.
Table VII: Summary of Open Research Challenges in HAPS-Enabled TN/NTN Integration
Research Challenge (Time Horizon, Type) Key Open Issues
Channel Modeling (Short-term, Technical) Especially at mmWave and higher frequencies, joint consistent modeling spanning terrestrial–HAPS–satellite links under non-stationarity, large coverage, and weather-/terrain-dependent effects.
Spectrum Management (Short- to Long-term, Technical/Regulatory) Interference cooperation and scalable spectrum sharing across diverse tiers with dense deployments and dynamic beams.
Mobility Management and Handover (Short-term, Technical) Robust signaling and precise handover time in the face of diverse propagation delays, quasi-stationary platforms, and large coverage areas.
Multi-Connectivity (Short-term, Technical) Bearer splitting/duplication across heterogeneous networks under dynamic mobility: control-plane overhead, synchronization, and scalability.
Interoperability and Standardization (Long-term, Technical/Standardization) Protocol development beyond existing norms to facilitate real-time multi-tier coordination, huge connection, and large delays.
Energy Efficiency (Short- to Long-term, Technical) Energy harvesting, storage, and energy-conscious scheduling are all components of sustainable HAPS operation under stringent energy limitations.
Legal and Regulatory Issues (Long-term, Regulatory) Large-scale HAPS deployment requires international cooperation, safety, spectrum control, and airspace management.
Deployment Costs (Long-term, Economic) For widespread adoption, public-private collaborations, sustainable business models, and economic viability are required.
Security, Jamming, and Adversarial ML Threats (Short- to Long-term, Technical) Protection in wide-area LoS situations from hostile ML attacks, spoofing, cyberthreats, and jamming.
Channel Modeling: As previously examined in the literature (e.g., [251, 97, 237, 170, 241]), Doppler effects, delay spread, and atmospheric attenuation are among the important propagation characteristics of HAPS-enabled NTNs that have been examined under a variety of assumptions. However, the requirement for jointly consistent representations across diverse links and large geographic areas keeps channel modeling open in HAPS-enabled TN/NTN integration. Channel coherence is affected by non-stationarity brought on by platform motion, altitude changes, and atmospheric dynamics, while spatial heterogeneity is introduced by effects that rely on the topography and the weather. At mmWave and higher frequencies, when attenuation and beam misalignment can significantly impair link reliability, these difficulties are made worse.
Spectrum: As reported in earlier studies on spectrum sharing and interference management (e.g., [80, 140, 227, 134]), spectrum coexistence between HAPS, satellite, and terrestrial systems is a well-known challenge. This problem is exacerbated by diverse coverage scales, dynamic beamforming, and dense node distributions in HAPS-enabled multi-tier deployments. Scalable coordination methods, scheduling that considers interference, and adaptive spectrum access techniques that function well in the face of mobility and quickly shifting interference situations are among the unresolved difficulties.
Multi-connectivity: It has been demonstrated that multi-connectivity can support strict URLLC requirements, including in circumstances related to NTN [187, 186]. But even with the Stage-2 architecture outlined in 3GPP TS 37.340 [7], there is still room for multi-connectivity advantages to be practically realized in HAPS-enabled NTN deployments. Syncing across heterogeneous links, scalability under dynamic mobility situations, and higher signaling overhead are among the main obstacles [183].
Mobility Management: Numerous studies have been conducted on mobility and handover in HAPS-assisted NTNs [112, 113, 131]. However, mobility management is still a challenge in HAPS-enabled TN/NTN integration because of diverse propagation delays, quasi-stationary aerial platforms, and vast coverage regions. The terrestrial assumptions that underlie current LTE/NR mobility procedures [6, 9] may result in inefficient handover timing and signaling, especially for services that are sensitive to latency and reliability under high mobility and limited signal strength gradients [240].
Interoperability: Interoperability is still a major obstacle, as stated in current TN/NTN standardization initiatives (e.g., [2, 4]). Although the present standards offer a starting point, more protocol development is needed to accommodate huge connectivity, substantial propagation delays, and real-time coordination between satellites, HAPS, and terrestrial nodes.
Energy: Energy efficiency is still a significant challenge, as was previously mentioned in the context of energy-efficient HAPS operation (e.g., [177, 41]). In order to guarantee continued operation under changing lighting and traffic situations, there are still unresolved challenges with effective energy harvesting, long-term storage, dynamic power allocation, and energy-aware scheduling.
Legal Regulations: Large-scale HAPS deployment presents unresolved issues with airspace management, safety, spectrum regulation, and international cooperation, despite the fact that regulatory features of aerial and non-terrestrial platforms have been taken into account in previous work.
Deployment Costs: The economic viability of large-scale HAPS deployment is still up for debate, despite the fact that techno-economic elements have been examined in previous research. To justify long-term investment, viable revenue models and public-private partnerships are necessary due to high deployment and maintenance costs.
Security, Jamming, and Adversarial Machine Learning Threats: As was previously mentioned in the context of HAPS-enabled NTN security [51, 242, 179, 230, 234, 239, 81, 205, 42, 169, 95, 228], HAPS-enabled systems are susceptible to cyberattacks, spoofing, and jamming due to their exposed wireless configuration. Additionally, the growing dependence on machine learning for optimization and control raises the possibility of adversarial attacks, underscoring the necessity of strong and reliable security measures.
VIII Discussion and conclusion